OpenStudy (anonymous):

which of the following is the equation of the line passing through the points c(-1,1) and D(7,-1)

OpenStudy (anonymous):

a. y+1=1/4(x-1) b. y-1=-1/4(x+1) c. y-1=4(x+1) d. y-1=1/4(x+1)

OpenStudy (anonymous):

is it just line passing or finding the slope to

OpenStudy (anonymous):

just the line

OpenStudy (anonymous):

okay, let me work on it

OpenStudy (anonymous):

ok (:

OpenStudy (rajat97):

use the formula of the line in two poin form (y-y1)/(y1-y2)=(x-x1)/(x1-x2) let any point be (x1,y1) so the other will become (x2,y2) and by putting the values in the equation , you will get the equation of the line as b)y-1=-1/4(x+1)

OpenStudy (anonymous):

i did it and i did it in to different ways and got two different answers i got D y-1=1/4(x+1) and A y+1=1/4(x-1)

OpenStudy (anonymous):

sorry if that doesn't help